Advertisements
Advertisements
Question
Forty percent of business travellers carry a laptop. In a sample of 15 business travelers, what is the probability that atleast three of the travelers have a laptop?
Advertisements
Solution
Given n = 5
p = `40/10 = 2/5`
q = 1 – p = `1 - 2/5 = (5-2)/5 = 3/5`
The binomial distribution P(X = x) = `15"c"_x (2/5)^x (3/5)^(15 - x)`
P(at least three of the travelers have a laptop)
= P(X ≥ 3)
P(X ≥ 3) = 1 – P(X < 3)
= 1 – [P(X = 0) + P(X = 1) + P(X = 2)]
P(X < 3) = P(X = 0) + P(X = 1) + P(X = 2)
= `15"c"_0 (2/5)^0 (3/5)^(15 - 0) + 15"c"_1 (2/5)^1 (3/5)^(15 - 1) + 15"c"_2 (2/5)^2 (3/5)^(15 - 2)`
= `(1)(1)(3/5)^15 + 15(2/5)(3/5)^14 + 105(2/5)^2(3/5)^13`
= `13^15/5^15 +(15 xx 2 xx 3^14)/5^5 +(105 xx 4 xx 3^13)/5^15`
= `(3^13 [3^2 + (30 xx 3) + (105 xx 4 xx 9)])/5^15`
= `(1.593 xx 10^6 xx (9 + 90 + 3780))/(3.055 xx 10^10)`
= `(1.593 xx 3879)/(3.055 xx 10^4)`
= `6179.247/30550`
= 0.20226
∴ P(X ≥ 3) = `1 - 020226`
= 0.79774
APPEARS IN
RELATED QUESTIONS
Write down the condition for which the binomial distribution can be used.
Among 28 professors of a certain department, 18 drive foreign cars and 10 drive local made cars. If 5 of these professors are selected at random, what is the probability that atleast 3 of them drive foreign cars?
Out of 750 families with 4 children each, how many families would be expected to have atmost 2 girls
Forty percent of business travellers carry a laptop. In a sample of 15 business travelers, what is the probability that 3 will have a laptop?
An experiment succeeds twice as often as it fails, what is the probability that in next five trials there will be at least three successes
Write the conditions for which the poisson distribution is a limiting case of binomial distribution
The average number of phone calls per minute into the switchboard of a company between 10.00 am and 2.30 pm is 2.5. Find the probability that during one particular minute there will be exactly 3 calls
The distribution of the number of road accidents per day in a city is poisson with mean 4. Find the number of days out of 100 days when there will be at most 3 accidents
Choose the correct alternative:
In turning out certain toys in a manufacturing company, the average number of defectives is 1%. The probability that the sample of 100 toys there will be 3 defectives is
Choose the correct alternative:
An experiment succeeds twice as often as it fails. The chance that in the next six trials, there shall be at least four successes is
